Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.30 pounds $ 16.60 per pound $ 38.18
Direct labor 1.00 hours $ 15.70 per hour $ 15.70
Variable manufacturing overhead 1.00 hours $ 9.30 per hour $ 9.30
Total standard cost per unit $ 63.18
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 572,700 $ 10,585 F $ 33,200 U
Direct labor $ 235,500 $ 3,200 U $ 15,700 U
Variable manufacturing overhead $ 139,500 $ 4,900 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many pounds of direct material were purchased and used in production?

2. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

3. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

Actual units = 572700/38.18 = 15000 Units

1) Material quantity variance = (Standard quantity-actual quantity)Standard rate

-33200 = (15000*2.3*16.60-16.60X)

16.60X = 572700+33200

Actual pound of material purchased and used = 36500 Pounds

2) Material price variance = (Standard price-actual price)actual quantity

10585 = (16.60*36500-36500X)

36500X = 605900-10585

X(actual cost per pound) = 16.31 per pound

3) labor efficiency variance= (Standard hour-actual hour)Standard rate

-15700 = (15000*15.7-15.70X)

15.70X = 251200

X(actual hour) = 16000 hour

Variable overhead efficiency variance = (15000-16000)*9.3 = 9300 U

Actual variable overhead = 139500+4400 = $143900